Bringham Company issues bonds with a par value of $630,000 on their stated issue date. The bonds mature in 7 years and pay 8% annual interest in semiannual payments. On the issue date, the annual market rate for the bonds is 10%. (Table B.1, Table B.2, Table B.3, and Table B.4) (Use appropriate factor(s) from the tables provided.)

1. What is the amount of each semiannual interest payment for these bonds?
2. How many semiannual interest payments will be made on these bonds over their life?
3. Use the interest rates given to select whether the bonds are issued at par, at a discount, or at a premium.
4. Compute the price of the bonds as of their issue date.
5. Prepare the journal entry to record the bonds’ issuance.
Answer #1

1) Semi annual interest payment = Bonds Face Value*Interest rate*6/12

= $630,000*8%*6/12 = $25,200

Therefore the amount of each semiannual interest payment for these bonds is $25,200.

2) Semiannual interest payments over the bond life = Bond life in years*2 semiannual periods

= 7 yrs*2 semiannual period = 14 periods

Therefore total 14 semiannual interest payments will be made on these bonds over their life .

3) The annual market rate (i.e. 10%) is more than annual interest rate (i.e. 8%) which means that market price of bonds is more than par value of bonds. Therefore the bonds are issued at a premium.

4) Current Price of Bonds = PV of par value+PV of semi annual interest payments

Present Value of par value = Par Value*PVF(2.5%, 14)

= $630,000*0.70773 = $445,868

Present Value of Interest payments = Semi annual interest*PVAF(2.5%, 14)

= $25,200*11.69091 = $294,611

Current Price of Bonds = $445,868+$294,611 = $740,479

Therefore the price of the bonds as of their issue date is $740,479.

5) Journal Entries (Amounts in $)

No Account Titles and Explanations Debit Credit
1 Cash 740,479
Premium on Bonds Payable (740,479-630,000) 110,479
Bonds Payable 630,000
(To record the issue of bonds)